Does a 60” RCP culvert work for the 25 year flow rate design controls and prevent flooding of the critical structure given the following conditions: 200’ long culvert @0.30%. Invert elevation at tail end of pipe = 10.00. ODOT ½ headwall. Pavement elevation = 17.95. Critical Structure FFE = 18.65.

Q25 = 140 cfs, Q50 = 172 cfs, TW25 = 2.5’ TW50 = 2.9’
Answer #1

ANSWERReinforced concrete pipe Current Besign! A) 1) Eischarge : 295 = 140CFS 050 = 1490FS. Area of culvert = I xor A = 4602 A = 28@ L200 Om 60 pu! ! Discharge CFS 100 Hw depth. (ilo) Length Culvest = 2009 slope 0.9 100 -0.003 chart head for concrete pipeH-value=5 hos Tw or ((critical depth 0)) 200 50K-6-0-25 60 100+ V100 DETO) Head (4) ho=2.5 ho = (3.5+4) from cortical depth -4 (a) using inlet Control homograph. Holo 1.3 Hw=1-35% bom HwE 6.75 Ft (11) H-value = 4.5 Hwe 4.5+4.95- (900X0-003) HW-3.15 (